# MicroMoE: Fine-grained Load Balancing for Mixture-of-Experts with Token Scheduling

### v1

@@ 4. MicroMoE Overview
We introduce a novel expert parallelism strategy, namely MicroEP. MicroEP optimizes load balancing across GPUs within every micro-batch through token scheduling across EDP groups.
@@
Based on MicroEP, we propose MicroMoE, an efficient distributed MoE training system. Figure 4 illustrates the architecture of MicroMoE. Following PP, each device hosts multiple model layers. Each MoE layer consists of a token dispatcher and multiple local expert replicas. Expert replica placement is coordinated by a global placement manager residing on device 0. We briefly illustrate the workflow of MicroMoE as follows.
@@
Prerequisites. MicroEP obtains scheduling space from intersecting EDP groups, which require two conditions: the DP degree must exceed the EP degree, and each GPU must host multiple expert replicas. We find that these requirements are satisfied in many open-sourced MoE models (Pangu; DeepSeek-V3). For example, DeepSeek-V3’s pre-training configuration employs 64-way EP and 128-way DP, with each GPU hosting 8 experts (DeepSeek-V3).
@@
Configuration. MicroEP introduces an integer parameter $d$, where $1<d\leq\frac{\text{DP\_degree}}{\text{EP\_degree}}$. MicroEP merges every $d$ EP groups into a MicroEP group. In the following sections, we consider the simplest case where $d=\frac{\text{DP\_degree}}{\text{EP\_degree}}$, i.e., the entire DP group forms a single MicroEP group.
@@
Initialization. During model initialization, the placement manager generates expert placements within all MicroEP groups and broadcasts the placements to all devices. The placement strategies are detailed in §6.
@@
Runtime. During model training, MicroEP balances GPU workloads by scheduling tokens within the MicroEP group. Specifically, after the gate networks assign tokens to specific experts, the dispatchers route tokens to particular expert replicas within their respective EDP groups. The scheduling algorithm is detailed in §5.
@@
After scheduling, the dispatchers initiate an all-to-all (dispatch) operation within the MicroEP group, sending tokens to their routed expert replicas. Then, each GPU executes FFN computation to process received tokens using its local expert replicas. Afterward, a second all-to-all (combine) operation returns processed tokens to their source GPUs. These operations follow the same patterns as vanilla EP, but occur in a communication group $d$ times larger.
@@
An example. Surprisingly, MicroEP can almost always achieve perfect load balance among GPUs, unless the expert load distribution is extremely skewed (as shown in §7.3). We can intuitively learn how MicroEP achieves such good performance from the example in Figure 3(c). Since expert 3 is heavily loaded, it places 8 tokens on GPU 3, preventing expert 0 from placing any token on the same GPU. Fortunately, expert 0 can place all its 4 tokens on GPU 0. Similarly, expert 2 also “shifts” some load from GPU 2 to GPU 1, while expert 1 shifts some load from GPU 1 to GPU 0. Consequently, the computational burden of heavily loaded experts is effectively distributed across the entire MicroEP group.
5. Token Scheduling
@@
Token scheduling is the core of MicroEP for fine-grained load balancing. The scheduling algorithm runs in two steps: The first step is to distribute expert loads among their replicas, balancing GPU loads (§5.1). The second step is to route tokens to specific expert replicas, enforcing the calculated replica loads (§5.2). Then, we discuss where and when to launch the scheduling in §5.3 and §5.4. Finally, we discuss some additional scheduling optimizations in Appendix A.
5.1. Determining Replica Loads
@@
Table 1. List of Symbols and Notations.
@@
$G_{MicroEP}$ set of GPUs in the concerned MicroEP group
$d$ a parameter $\frac{|G_{MicroEP}|}{\text{EP\_degree}}$
@@
  • Notations. Let $E$ represent the set of all experts, $G_{MicroEP}$ represent the set of GPUs in the concerned MicroEP group, and $G_{EDP}^{e}$ represent the set of GPUs in the EDP group of expert $e$. Since we currently focus on a single MicroEP group, we assume that $G_{EDP}^{e}\subseteq G_{MicroEP},\forall e\in E$.
@@
  • Constraints. Let $load_{e}(e\in E)$ denote the total load (number of tokens) of expert $e$ in the concerned MicroEP group. A valid distribution must ensure that each expert distributes its total load across its replicas.
  • Objective. Due to synchronization requirements for all-to-all communication before and after expert computation, the GPU with the highest load becomes the performance bottleneck. Therefore, we should minimize the maximum GPU load across the MicroEP group.
@@
(1) minimize $\max\limits_{g\in G_{MicroEP}}\left\{\sum\limits_{e\in E:g\in G_{EDP}^{e}}x_{e}^{g}\right\},$
@@
Problem 1 is a linear programming problem (LPP), which can be efficiently solved in polynomial time relative to the number of GPUs and experts. The number of variables is $O(|E|d)$, and the number of constraints is $O(|E|+|G_{MicroEP}|)$. Given its modest scale, we solve this LPP using a single CPU thread with the HiGHs solver (HiGHS). GPU acceleration or multi-threading would not yield significant performance benefits for this scale of optimization.
@@
5.2. Routing Tokens to Replicas
@@
To enhance efficiency, we can manipulate token ranges rather than individual tokens. Let $input_{e}^{g}$ denote the input load of expert $e$ from GPU $g$, i.e., the number of tokens on GPU $g$ assigned to expert $e$ ($\sum_{g\in G_{MicroEP}}input_{e}^{g}=load_{e}$). The pseudo code of token routing is shown in Algorithm 1, Lines 10-16.
Locality-aware routing. Previous research has underscored that all-to-all communication introduces significant overhead to MoE systems (FasterMoE; Tutel; Lina). To reduce the all-to-all communication volume, we can leverage data locality during token routing. Specifically, when GPU $g$ holds a replica of expert $e$, we prioritize routing tokens from GPU $g$ to its local expert replica before considering remote replicas. The pseudo code is shown in Algorithm 1, Lines 4-9. Furthermore, we can consider communication overhead during optimization, which is detailed in Appendix A.1.
@@ 17       $remain\_x_{e}^{g^{\prime}}\leftarrow remain\_x_{e}^{g^{\prime}}-y$
5.3. Distributed Scheduling across Devices
As described in §5.1 and §5.2, the scheduling algorithm requires the global load information ($input_{e}^{g}$) across the entire MicroEP group as input. This raises the question of where to place the scheduler.
We consider two candidate locations for the scheduler: centralized on one device or distributed across all devices. A centralized scheduler needs to gather load information from all devices, perform the scheduling, and scatter the results. Alternatively, distributed schedulers require all devices to perform an all-gather operation to collect global load information and execute the scheduling algorithm independently.
@@ 5.3. Distributed Scheduling across Devices
Notably, the distributed scheduling maintains consistency because MicroEP’s scheduling algorithm is deterministic—all devices use identical inputs, follow identical algorithms, and produce identical results.
5.4. Overlapping Scheduling to Hide Latency
@@
As described in §4, the scheduling executes immediately after the gate network and before the all-to-all communication. We observe that existing distributed training frameworks usually perform some operations during this period. For instance, Megatron-LM (Megatron-LM) executes a token permutation operation to replicate tokens by top-K times and sort tokens by expert indices before all-to-all dispatching. Therefore, we can overlap the scheduling on CPUs with the permutation on GPUs. For frameworks without suitable overlapping operations, we propose a pipeline mechanism to hide scheduling latency, which is detailed in Appendix A.2.
6. Expert Placement
Expert placement is fundamental to the load balancing performance of token scheduling in MicroEP. As shown in LPP 1, the expert placement determines the EDP groups ($G_{EDP}^{e}$), which are the key components of the constraint matrix and thus significantly affect the optimization result. Although our experiments demonstrate that some simple placements, such as random shuffling, usually provide satisfactory results (§7.3), identifying optimal placement strategies is still crucial for maximizing system performance. In the following sections, we first analyze “What is an optimal expert placement?” in §6.1 and then discuss “How to construct an optimal expert placement?” in §6.2 and §6.3. Finally, we propose an adaptive replacement mechanism in §6.4.
6.1. Analysis of Optimal Expert Placement
Figure 5. The graph abstraction of an example expert placement. Color bars indicate edges (experts) between vertices (GPUs). $G_{max}$ contains GPU 0, 3. Expert 0 is entirely in $G_{max}$. Experts 1,3 partially intersect with $G_{max}$ and cannot distribute any load within $G_{max}$.
@@ 6.1. Analysis of Optimal Expert Placement
Proof.
Among all optimal solutions, we choose the solution with the minimal $|G_{max}|$ (number of GPUs with load $m$). Consider experts whose EDP groups partially intersect with $G_{max}$. Suppose that the EDP group of expert $e$ consists of two GPUs $g_{0}\in G_{max}$ and $g_{1}\notin G_{max}$. If expert $e$ did distribute any load on GPU $g_{0}$, it could move a tiny amount of load from $g_{0}$ to $g_{1}$. Afterward, the loads of both $g_{0}$ and $g_{1}$ would be less than $m$, yielding a new optimal solution with smaller $|G_{max}|$. The new solution would contradict our choice of the solution with the minimal $|G_{max}|$. Therefore, expert $e$ cannot distribute any load on GPU $g_{0}$. ∎
@@
(2) $m\cdot|G_{max}|=\sum\limits_{e\in E:G_{EDP}^{e}\subseteq G_{max}}load_{e}$
@@
(3) $m=\max_{G_{max}\subseteq G_{MicroEP}}\left\{\frac{1}{|G_{max}|}\sum\limits_{e\in E:G_{EDP}^{e}\subseteq G_{max}}load_{e}\right\}$
@@
Graph abstraction of expert placement. Furthermore, we formulate expert placement with graph theory. Let each GPU $g\in G_{MicroEP}$ represent a vertex, and each expert $e\in E$ represent a hyperedge connecting all GPUs in $G_{EDP}^{e}$. Hence, an expert placement can be represented by an undirected hypergraph [^2] When $d=2$, the hypergraph is a conventional graph. For simplicity, we omit the prefix “hyper” in the rest of this paper. , denoted by $\mathcal{G}(G_{MicroEP},\{G_{EDP}^{e}:e\in E\})$. Figure 5 shows an example of the graph and the optimal solution corresponding to an expert placement.
@@
6.2. Symmetric Placement without Real Expert Loads
If we have no prior knowledge of the real expert load distribution, we can construct symmetric placements, treating all experts equally. In such scenarios, we can assume that all expert loads follow an independent and identically distributed (i.i.d.) pattern. Thereby, the problem becomes: Given the number of vertices and edges, and the identical distribution of edge weights, how to construct a graph that minimizes the expectation of the maximum induced subgraph density?
We recognize the challenge of the above problem due to the vast space of possible graphs and distributions. Nevertheless, we propose a near-optimal symmetric placement strategy for many practical configurations using Cayley graphs (cayley). Our intuition is that the inherent symmetry of Cayley graphs ensures a balanced distribution of edges across vertices, preventing some induced subgraphs from having significantly larger density than others. Since Cayley graphs involve complex group theory, we illustrate our construction method in Appendix B. We demonstrate the effectiveness of our constructions through both example studies (Appendix B.2) and experimental evaluations (§7.3).
6.3. Asymmetric Placement with Real Expert Loads
If we know real expert load distributions in advance, we can construct asymmetric placements based on specific distributions. Unlike symmetric scenarios, we can vary both replica counts and replica locations across different experts for better load balancing, similar to previous works (SmartMoE; FlexMoE; Prophet).
@@ 6.3. Asymmetric Placement with Real Expert Loads
6.4. Adaptive Replacement
Based on the symmetric and asymmetric placements, we further propose an adaptive replacement (AR) mechanism for MicroMoE to optimize performance under dynamic expert loads.
@@ 6.4. Adaptive Replacement
Specifically, for moderately imbalanced workloads, token scheduling effectively maintains optimal balance with static, symmetric placements (as shown in §7.3). For highly skewed workloads, MicroMoE adopts asymmetric placements to mitigate coarse-grained imbalances before using token scheduling for fine-grained optimization. Since asymmetric placements require real-time expert loads, MicroMoE adopts adaptive replacement to monitor expert load distributions and adjust placements when significant distributional shifts are detected.
Implementation of adaptive replacement. We implement the adaptive replacement mechanism in MicroMoE using the placement manager (according to Figure 4). (1) During model initialization, the placement manager initializes the model states of all devices using the symmetric placement strategy, providing conservative and general load balancing capabilities. (2) During training, the placement manager monitors expert load information within each micro-batch in the background. (3) For every few iterations, the placement manager predicts future load distributions using historical data with time series analysis techniques, such as moving averages (Prediction). Then, it evaluates the performance of current placements on future distributions using Equation 3. If the future performance drops below a specific threshold, the placement manager generates new optimal asymmetric placements and re-initializes global model states accordingly.
Difference between MicroMoE’s adaptive replacement and expert scheduling solutions. The system implementation of MicroMoE’s adaptive replacement is similar to existing expert scheduling solutions, such as FlexMoE (FlexMoE) and SmartMoE (SmartMoE). Nevertheless, their design goals and algorithms are fundamentally different:
Design goals: In FlexMoE, changing expert placement is the only means for load balancing. However, in MicroMoE, the primary weapon is token scheduling, while adaptive replacement is a further optimization to token scheduling. Even with static placement, MicroMoE can still achieve good load balancing performance at the micro-batch granularity.
Algorithms: The different design goals of FlexMoE and MicroMoE lead to different placement algorithms. In FlexMoE, all replicas of an expert have identical loads, i.e., the replica load is simply calculated as the expert load divided by the replica count. In contrast, the replica loads in MicroMoE are determined by the LPP 1. Therefore, MicroMoE requires the graph theory in §6.1 to guide the placement strategy.
7. Evaluation
7.1. Experimental Setup

### v2

@@ 4 FineMoE Overview
We introduce a novel expert parallelism strategy, namely FineEP. FineEP optimizes load balancing across GPUs within every micro-batch through token scheduling across EDP groups.
@@
Based on FineEP, we propose FineMoE, an efficient distributed MoE training system. Figure 4 illustrates the architecture of FineMoE. Following PP, each device hosts multiple model layers. Each MoE layer consists of a token dispatcher and multiple local expert replicas. Expert replica placement is coordinated by a global placement manager residing on device 0. We briefly illustrate the workflow of FineMoE as follows.
@@
Prerequisites. FineEP obtains scheduling space from intersecting EDP groups, which require two conditions: the DP degree must exceed the EP degree, and each GPU must host multiple expert replicas. We find that these requirements are satisfied in many open-sourced MoE models [Pangu, DeepSeek-V3]. For example, DeepSeek-V3’s pre-training configuration employs 64-way EP and 128-way DP, with each GPU hosting 8 experts [DeepSeek-V3].
@@
Configuration. FineEP introduces an integer parameter $d$, where $1<d\leq\frac{\text{DP\_degree}}{\text{EP\_degree}}$. FineEP merges every $d$ EP groups into a FineEP group. In the following sections, we consider the simplest case where $d=\frac{\text{DP\_degree}}{\text{EP\_degree}}$, i.e., the entire DP group forms a single FineEP group.
@@
Initialization. During model initialization, the placement manager generates expert placements within all FineEP groups and broadcasts the placements to all devices. The placement strategies are detailed in §6.
@@
Runtime. During model training, FineEP balances GPU workloads by scheduling tokens within the FineEP group. Specifically, after the gate networks assign tokens to specific experts, the dispatchers route tokens to particular expert replicas within their respective EDP groups. The scheduling algorithm is detailed in §5.
@@
After scheduling, the dispatchers initiate an all-to-all (dispatch) operation within the FineEP group, sending tokens to their routed expert replicas. Then, each GPU executes FFN computation to process received tokens using its local expert replicas. Afterward, a second all-to-all (combine) operation returns processed tokens to their source GPUs. These operations follow the same patterns as vanilla EP, but occur in a communication group $d$ times larger.
@@
An example. Surprisingly, FineEP can almost always achieve complete load balance among GPUs, unless the expert load distribution is extremely skewed (as shown in §7.3). We can intuitively learn how FineEP achieves such good performance from the example in Figure 3(c). Since expert 3 is heavily loaded, it places 8 tokens on GPU 3, preventing expert 0 from placing any token on the same GPU. Fortunately, expert 0 can place all its 4 tokens on GPU 0. Similarly, expert 2 also “shifts” some load from GPU 2 to GPU 1, while expert 1 shifts some load from GPU 1 to GPU 0. Consequently, the computational burden of heavily loaded experts is effectively distributed across the entire FineEP group.
5 Token Scheduling
@@
Token scheduling is the core of FineEP for fine-grained load balancing. The scheduling algorithm runs in two steps: The first step is to distribute expert loads among their replicas, balancing GPU loads (§5.1). The second step is to route tokens to specific expert replicas, enforcing the calculated replica loads (§5.2). Then, we discuss where and when to launch the scheduling in §5.3 and §5.4. Finally, we discuss some additional scheduling optimizations in Appendix A.
5.1 Determining Replica Loads
@@
Table 1: List of Symbols and Notations.
@@
$G_{FineEP}$ set of GPUs in the concerned FineEP group
$d$ a parameter $\frac{|G_{FineEP}|}{\text{EP\_degree}}$
@@
  • Notations. Let $E$ represent the set of all experts, $G_{FineEP}$ represent the set of GPUs in the concerned FineEP group, and $G_{EDP}^{e}$ represent the set of GPUs in the EDP group of expert $e$. Since we currently focus on a single FineEP group, we assume that $G_{EDP}^{e}\subseteq G_{FineEP},\forall e\in E$.
@@
  • Constraints. Let $load_{e}(e\in E)$ denote the total load (number of tokens) of expert $e$ in the concerned FineEP group. A valid distribution must ensure that each expert distributes its total load across its replicas.
  • Objective. Due to synchronization requirements for all-to-all communication before and after expert computation, the GPU with the highest load becomes the performance bottleneck. Therefore, we should minimize the maximum GPU load across the FineEP group.
@@
minimize $\max\limits_{g\in G_{FineEP}}\left\{\sum\limits_{e\in E:g\in G_{EDP}^{e}}x_{e}^{g}\right\},$ (1)
@@
Problem 1 is a linear programming problem (LPP), which can be efficiently solved in polynomial time relative to the number of GPUs and experts. The number of variables is $O(|E|d)$, and the number of constraints is $O(|E|+|G_{FineEP}|)$. Given its modest scale, we solve this LPP using a single CPU thread with the HiGHs solver [HiGHS]. GPU acceleration or multi-threading would not yield significant performance benefits for this scale of optimization.
@@
5.2 Routing Tokens to Replicas
@@
To enhance efficiency, we can manipulate token ranges rather than individual tokens. Let $input_{e}^{g}$ denote the input load of expert $e$ from GPU $g$, i.e., the number of tokens on GPU $g$ assigned to expert $e$ ($\sum_{g\in G_{FineEP}}input_{e}^{g}=load_{e}$). The pseudo code of token routing is shown in Algorithm 1, Lines 10-16.
Locality-aware routing. Previous research has underscored that all-to-all communication introduces significant overhead to MoE systems [FasterMoE, Tutel, Lina]. To reduce the all-to-all communication volume, we can leverage data locality during token routing. Specifically, when GPU $g$ holds a replica of expert $e$, we prioritize routing tokens from GPU $g$ to its local expert replica before considering remote replicas. The pseudo code is shown in Algorithm 1, Lines 4-9.
Communication-aware scheduling. Furthermore, we can consider communication overhead during scheduling. Specifically, we can reformulate the optimization problem to minimize the overall maximum execution time, incorporating both computation and communication. Moreover, we can model the heterogeneity between intra-node and inter-node communication. Due to space limits, the modified optimization problem is detailed in Appendix A.1.
@@ 18       $remain\_x_{e}^{g^{\prime}}\leftarrow remain\_x_{e}^{g^{\prime}}-y$
5.3 Distributed Scheduling across Devices
As described in §5.1 and §5.2, the scheduling algorithm requires global load information ($input_{e}^{g}$) across the entire FineEP group to generate a token dispatching plan. This raises the question of where to place the scheduler.
We consider two candidate locations for the scheduler: centralized on one device or distributed across all devices. A centralized scheduler needs to gather load information from all devices, perform the scheduling, and scatter the results. Alternatively, distributed schedulers require all devices to perform an all-gather operation to collect global load information and execute the scheduling algorithm independently.  [^1] The distributed scheduling approach maintains consistency because FineEP’s scheduling algorithm is deterministic.
@@ 5.3 Distributed Scheduling across Devices
5.4 Overlapping Scheduling to Hide Latency
@@
As described in §4, the scheduling executes immediately after the gate network and before the all-to-all communication. We observe that existing distributed training frameworks usually perform some operations during this period. For instance, Megatron-LM [Megatron-LM] executes a token permutation operation to replicate tokens by top-K times and sort tokens by expert indices before all-to-all dispatching. Therefore, we can overlap the scheduling on CPUs with the permutation on GPUs. For frameworks without suitable overlapping operations, we propose a pipelining mechanism to hide scheduling latency, which is detailed in Appendix A.2.
6 Expert Placement
Expert placement is fundamental to the load balancing performance of token scheduling in FineEP. Figure 6(a) shows the relationship between expert placement and token scheduling in MicroMoE. Specifically, according to LPP 1, the expert placement determines the EDP groups ($G_{EDP}^{e}$), which are the key components of the constraint matrix and thus significantly affect the optimization result. Although our experiments demonstrate that some simple placement methods, such as random shuffling, usually provide acceptable results (§7.3), identifying optimal placement strategies is still crucial for maximizing system performance. In the following sections, we first analyze “What is an optimal expert placement?” in §6.1 and then discuss “How to construct an optimal expert placement?” in §6.2 and §6.3. Finally, we propose an adaptive replacement mechanism in §6.4.
6.1 Analysis of Optimal Expert Placement
Figure 5: The graph abstraction of an example expert placement. Color bars indicate edges (experts) between vertices (GPUs). $G_{max}$ contains GPU 0, 3. Expert 0 is entirely in $G_{max}$. Experts 1,3 partially intersect with $G_{max}$ and cannot distribute any load within $G_{max}$.
@@ 6.1 Analysis of Optimal Expert Placement
[…]
@@
$m\cdot|G_{max}|=\sum\limits_{e\in E:G_{EDP}^{e}\subseteq G_{max}}load_{e}$ (2)
@@
$m=\max_{G_{max}\subseteq G_{FineEP}}\left\{\frac{1}{|G_{max}|}\sum\limits_{e\in E:G_{EDP}^{e}\subseteq G_{max}}load_{e}\right\}$ (3)
@@
Graph abstraction of expert placement. Furthermore, we formulate expert placement with graph theory. Let each GPU $g\in G_{FineEP}$ represent a vertex, and each expert $e\in E$ represent a hyperedge connecting all GPUs in $G_{EDP}^{e}$. Hence, an expert placement can be represented by an undirected hypergraph denoted by $\mathcal{G}(G_{FineEP},\{G_{EDP}^{e}:e\in E\})$. (When $d=2$, the hypergraph is a conventional graph. For simplicity, we omit the prefix “hyper” in the rest of this paper.) Figure 5 shows an example of the graph and the optimal solution corresponding to an expert placement.
@@
6.2 Symmetric Placement without Expert Loads
If we have no prior knowledge of the real expert load distribution, we can construct symmetric placements, treating all experts equally. Symmetric placements provide conservative and general load balancing capability in terms of unknown load distributions. In such scenarios, we can assume that all expert loads follow an independent and identically distributed (i.i.d.) pattern. Thereby, the problem becomes: Given the number of vertices and edges as well as the identical distribution of edge weights, how to construct a graph that minimizes the expectation of the maximum induced subgraph density?
We recognize the challenge of the above problem due to the vast space of possible graphs and distributions. Nevertheless, we propose a near-optimal symmetric placement strategy for many practical configurations using Cayley graphs [cayley]. Our intuition is that the inherent symmetry of Cayley graphs ensures a balanced distribution of edges across vertices, preventing some induced subgraphs from having significantly larger density than others. Since Cayley graphs involve complex group theory, we illustrate our construction method in Appendix B.
6.3 Asymmetric Placement with Expert Loads
If we know real expert load distributions in advance, we can construct asymmetric placements tailored to them. Unlike symmetric scenarios, we can vary both replica counts and replica locations across different experts for better load balancing, similar to previous works [SmartMoE, FlexMoE, Prophet].
@@ 6.3 Asymmetric Placement with Expert Loads
(a) Data flow and analysis flow of MicroMoE.
(b) Data flow of expert scheduling systems.
Figure 6: Differences between MicroMoE and expert scheduling systems.
6.4 Adaptive Replacement
Based on symmetric and asymmetric placements, we further propose an adaptive replacement (AR) mechanism for FineMoE to optimize performance under dynamic expert loads.
@@ 6.4 Adaptive Replacement
Specifically, for moderately imbalanced workloads, token scheduling sufficiently maintains complete balance with static, symmetric placements (as shown in §7.3). For highly skewed workloads, FineMoE adopts asymmetric placements to mitigate coarse-grained imbalances before using token scheduling for fine-grained optimization. Since asymmetric placements require real-time expert loads, FineMoE adopts adaptive replacement to monitor expert load distributions and adjust placements when significant distributional shifts are detected.
Implementation of adaptive replacement. We implement the adaptive replacement mechanism in FineMoE using the placement manager (according to Figure 4). (1) During model initialization, the placement manager initializes the model states of all devices using the symmetric placement strategy, providing conservative and general load balancing capabilities. (2) During training, the placement manager monitors expert load information within each micro-batch in the background. (3) For every few iterations, the placement manager predicts future load distributions using historical data with time series analysis techniques, such as moving averages [Prediction]. Then, it evaluates the performance of current placements on future distributions using Equation 3. If the future performance drops below a specific threshold, the placement manager generates new optimal asymmetric placements and re-initializes global model states accordingly.
Difference between FineMoE’s adaptive replacement and expert scheduling solutions. The system implementation of FineMoE’s adaptive replacement is similar to existing expert scheduling solutions, such as FlexMoE [FlexMoE] and SmartMoE [SmartMoE]. Nevertheless, their design goals and algorithms are fundamentally different:
Design goals: In systems like FlexMoE, changing expert placement is the only means for load balancing, as shown in Figure 6(b). However, in FineMoE, the primary weapon is token scheduling, while adaptive replacement is a further optimization to token scheduling, as shown in Figure 6(a). Even with static placement, FineMoE can still achieve good load balancing performance at micro-batch granularity.
Figure 7: End-to-end speedup of different systems compared with Megatron-LM.
Algorithms: Different design goals yield distinct placement algorithms. In existing expert scheduling systems, an expert’s replicas typically have identical loads (i.e., replica_load=expert_load/replica_count). Therefore, they can leverage greedy or dynamic programming algorithms accordingly [SmartMoE, FlexMoE]. In contrast, replica loads in FineMoE are determined by linear programming. Therefore, FineMoE requires the graph theory in §6.1 to guide the placement strategy.
7 Evaluation
7.1 Experimental Setup